OpenStudy (anonymous):

6x - 8 + 2x - 5 = 7 - 2x solve for x

Parth (parthkohli):

\( \color{Black}{\Rightarrow 8x - 13 = 7 - 2x}\) \( \color{Black}{\Rightarrow 8x = 20 - 2x}\) \( \color{Black}{\Rightarrow 8x + 2x = 20 }\) Can you manage from this point?

OpenStudy (anonymous):

10x=20

OpenStudy (anonymous):

x=2(:

mathslover (mathslover):

\[\huge{6x-8+2x-5=7-2x}\] \[\huge{6x+2x-8-5=7-2x}\] \[\huge{8x-13=7-2x}\] \[\huge{8x+2x=7+13}\] \[\huge{10x=20}\] \[\huge{x=\frac{20}{10}=2}\]